# Ambipolar doping of a charge-transfer insulator in the Emery model

### v1

@@ Abstract
Understanding the similarities and differences between adding or removing electrons from a charge-transfer insulator may provide insights about the origin of the electron-hole asymmetry found in cuprates. Here we study with cellular dynamical mean-field theory the Emery model set in the charge-transfer insulator regime, and dope it with either electrons or holes. We consider the normal state only and focus on the doping evolution of the orbital character of the dopants and on the nature of the doping driven transition. Regarding the orbital character of the dopants, we found an electron-hole asymmetry: doped electrons mostly enter the copper orbitals, whereas doped holes mostly enter the oxygen orbitals. Regarding the nature of the doping driven transition, we found no qualitative electron-hole asymmetry: upon either electron or hole doping, there is a two-stage transition from a charge-transfer insulator to a strongly correlated pseudogap and then to a metal. This shows that a strongly correlated pseudogap is an emergent feature of doped correlated insulators in two dimensions, in qualitative agreement with recent experiments on ambipolar Sr1-xLaxCuO2+y cuprate films. Our results indicate that merely doping with holes or electrons a charge-transfer insulator is not sufficient for explaining the electron-hole asymmetry observed in the normal state phase diagram of cuprates. Our work reinforces the view that actual hole-doped cuprates are more correlated than their electron-doped counterparts.
@@ I Introduction
Here, the problem we want to consider is what type of asymmetry results solely from doping a charge-transfer insulator in the normal state with either electrons or holes. First, is there any asymmetry in the doping evolution of the orbital character of the dopant carriers upon electron or hole doping? Second, is there any asymmetry in the qualitative features of the normal state doping driven transition upon electron or hole doping? To address these questions, we consider the finite-temperature normal state of the Emery model set in the charge-transfer insulating regime and dope the resulting charge-transfer insulator with either electron or holes (see sketch of the density of states of the undoped model in Fig. 1(a)).
The strong correlations giving rise to the charge-transfer insulator present an outstanding theoretical challenge, necessitating the use of nonperturbative methods. Over the years, significant progress on the physics of the Emery model has been enabled by advanced computational techniques, including quantum Monte Carlo methods [20, 21, 22, 23, 24, 25, 26, 27], density matrix renormalization group [28, 29, 30], two-particle self-consistent approach [31, 32], single-site dynamical mean-field theory [33, 34, 35, 36, 37, 8, 17, 38, 39] and its cluster extensions [40, 41, 42, 43, 44, 45, 46, 47, 48], variational cluster approximation [49, 50], density matrix embedding theory [51], and 2D tensor network methods [52]. However, due to the complexity of the Emery model, our understanding of the finite-temperature aspects of the normal state doping driven metal to charge-transfer insulator transition remain incomplete. Our work employs finite-temperature cellular extension [53, 54, 55] of dynamical mean-field theory [56] (CDMFT) and builds upon Ref. [44], which focused on hole doping only.
@@ II Summary of key findings
Second, we show that the low-temperature normal state doping driven metal to charge-transfer insulator transition has the same qualitative features upon either electron or hole doping: at zero doping, there is a continuous charge-transfer insulator to strongly correlated pseudogap transition, and at finite doping there is a discontinuous transition between the strongly correlated pseudogap and a more conventional correlated metal (see sequence of phases in Fig. 1(c)). The precise value of doping of the pseudogap to metal transition depends on the dopant carriers, but the qualitative features of the normal state doping driven transition of a charge-transfer insulator is thus symmetric upon electron or hole doping. This transition is analogous to that found in the doped Mott-Hubbard insulator realised with the single-band 2D Hubbard model [67, 68], as pointed out in Ref. [44] for hole doping.
Our findings that the ambipolar doping of a charge-transfer insulator (i) does produce an asymmetry in the orbital character of the dopants, (ii) but is not sufficient to produce an asymmetry on the qualitative features of the normal state doping driven transition, enable us to derive two main insights on the microscopic mechanism controlling the doping of a charge-transfer insulator, as we shall discuss in more detail in Section IX.
First, when ambipolar doping of a given charge-transfer insulating system is experimentally attainable, a qualitatively electron-hole symmetry in the normal state doping driven transition should be expected. Indeed, upon ambipolar doping of the cuprates films of SLCO [9], a surprising symmetry in the high-energy features of the bandstructure of the CuO2 planes is found. Most importantly, our results indicate that a strongly correlated pseudogap, namely a pseudogap solely arising from Mott physics plus short-range correlations, can emerge from a charge-transfer insulator upon either electron or hole doping, i.e. irrespective of the orbital character of the dopant carriers. This finding is in qualitative agreement with the work on ambipolar doping of cuprate films of SLCO [9]. Our finding implies that a strongly correlated pseudogap is an emergent feature of doped correlated insulators in two dimensions.
@@ III.1 Emery model
The Emery model takes into account three hoping processes. An electron can hop from an oxygen orbital (i) to a nearest-neighbor copper orbital via the hopping amplitude $t_{pd}$ (orange arrows), (ii) to a nearest-neighbor (diagonal) oxygen orbital via the hopping amplitude $t_{pp}$ (light blue arrows), and (iii) to an oxygen orbital across the copper orbital via the hopping amplitude $t_{pp}^{\prime}$ (dark blue arrows). The sign of the copper-oxygen and oxygen-oxygen hopping follows the convention indicated in Fig. 2. Let $\epsilon_{d}$ be the energy of the copper $3d_{x^{2}-y^{2}}$ orbitals and $\epsilon_{p}$ be the energy of the oxygen $2p_{x},2p_{y}$ orbitals. Regarding the interaction processes, on a copper orbital two electrons experience a Coulomb repulsion $U_{d}$. For simplicity, we neglect the Coulomb repulsion $U_{p}$ between electrons on an oxygen orbital.
@@ III.2 Cellular dynamical mean-field theory
To solve the Emery model Eq. 1 at finite temperature, we use the cellular extension [53, 54] of dynamical mean-field theory [56] (CDMFT) and follow the procedure described in Ref. [44]. Here we consider the CDMFT equations in the normal state only. The basic idea of CDMFT is to partition the lattice into a superlattice of clusters, to single out one cluster from the lattice and embed it in a self-consistent reservoir of noninteracting electrons. In this work we consider a cluster of $12$ sites with $N_{d}=4$ copper sites and $N_{p}=8$ oxygen sites, as indicated by the green square in Fig. 2.
@@ III.3 Model parameters of this work
Let us also comment our choice of $|\epsilon_{p}-\epsilon_{d}|=8$. This choice follows from the goal of our work to investigate the effect of doping a clear charge-transfer insulator. A smaller value of $|\epsilon_{p}-\epsilon_{d}|$ would be arguably closer to the values obtained in bandstructure calculations for cuprates [70, 8]. However, this would increase the mixed character between the $3d$ and $2p$ orbitals, which, in the Zaanen-Sawatzky-Allen framework [12], moves the Emery model towards the “intermediate” (or covalent or mixed-valence) region. In our work we wish to set the model in a clear charge-transfer insulator regime, that is, away from the intermediate regime. In the language of Ref. [48], our choice of parameters corresponds to an “ionic charge-transfer insulator”, and a smaller value of $|\epsilon_{p}-\epsilon_{d}|$ corresponds to a “covalent charge-transfer insulator”. Furthermore, small values of $|\epsilon_{p}-\epsilon_{d}|$ drastically reduce the Monte Carlo sign.
@@ IV.1 Two-stage metal to charge-transfer insulator transition
To summarise, upon either electron or hole doping, at low temperature the system undergoes a purely electronic two-stage doping driven metal insulator transition: a continuous charge-transfer insulator to pseudogap transition at $\delta=0$, and a discontinuous pseudpgap to correlated metal transition at finite $\delta$. This two-stage doping driven metal insulator transition also occurs in the doped Mott insulator realised by the single-band 2D Hubbard model solved with CDMFT [67, 68], as pointed out in Ref. [44].
IV.2 Peak in entropy and and enhanced charge compressibility
@@ V Temperature-doping phase diagram
Figure 5: Temperature $T$ vs doping $\delta$ normal state phase diagram of the Emery model solved with CDMFT. Data are for the set of parameters of Eq. 19 and $U_{d}=11.1$. Left: electron doping, right: hole doping. Dark blue squares show the incompressible charge-transfer insulator (CTI) at $\delta=0$. The spinodal line marking the pseudogap (PG) boundary is shown by open red triangles. The spinodal line marking the correlated metal (CM) boundary is shown by open orange triangles. The spinodal lines eventually merge at a critical end point shown by a solid purple circle. The Widom line, determined from the locus of maximum charge compressibility $\kappa$ vs $\delta$ in Fig. 4, extends from the critical end point and is indicated by a dashed purple line with open circles. The cyan dashed line with crosses shows the locus of maximum thermodynamic entropy $s$ vs $\delta$, as calculated by the crossing of the isotherms $\delta(\mu)$. These features occur on both the electron and hole-doped side of the phase diagram but are asymmetrical. Shaded grey area marks the inaccessible region to our calculations due to the Monte Carlo sign problem.
@@
The spinodal lines merge at a second order critical endpoint, for either hole or electron doping. The two filled purple circles are our estimates for the critical end points on the hole and electron doped side of the phase diagram. As a result, the pseudogap to metal transition is a purely electronic transition without symmetry breaking.
@@
Next, let us compare the hole and electron doping sides of the $T-\delta$ phase diagram. They appear qualitatively similar, but quantitatively different. In both cases, the system shows a first order transition between a pseudogap and a correlated metal which terminates in a critical end point at finite doping and finite temperature, and from which emerges a supercritical crossover characterised by enhanced density fluctuations. By changing the interaction strength $U_{d}$, in Section VIII we will show that the first order transition at hole doping and that at electron doping are connected in the $U_{d}-T-\delta$ phase diagram, thereby revealing their common origin in the Mott physics in two dimensions. In other words, in the $U_{d}-T-\delta$ diagram, there is a first-order surface transition.
For either electron or hole doping, the pseudogap to metal transition is followed at large doping by a maximum of the thermodynamic entropy $s(\delta)$. The locus of the maximum in $s(\delta)$ at different temperatures, as determined by the zero of the expansion coefficient $(dn_{\rm tot}/dT)_{\mu}$, is shown by a cyan dotted line with crosses: it is only weakly doping dependent and extends up to high temperatures. Physically, the peak in the thermodynamic entropy arises from the localization-delocalization physics of doped Mott insulators, as discussed in Ref. [75] in the context of the single-band 2D Hubbard model. These key features of the thermodynamic entropy $s$ – a doping-independent peak in $s(\delta)$ close to the value of doping where the pseudogap ends – are compatible with experimental results in hole-doped cuprates [78, 79].
Let us now turn to the quantitative differences between the two sides of the $T-\delta$ phase diagram. The temperature and doping position of the critical end point is asymmetric between electron and hole doping. The temperature of the critical end point is higher on the electron-doped side of the phase diagram. The origin of this asymmetry is not evident to us. A similar electron-hole asymmetry in the critical endpoint occurs in the single-band 2D Hubbard model on a triangular lattice [91]. This suggests that this asymmetry may be linked to the shape of the Fermi surface upon doping. However further investigations are needed to clarify this aspect.
@@
VI Phase characterization: local density of states
@@
The analysis in Sections IV and V reveals the existence of a first-order transition between two compressible metals at finite electron and hole doping. To further characterise the two metallic states separated by the transition and to show that they have the features of a strongly correlated pseudogap and of a correlated metal, we can turn to the study of the local density of states (DOS). This strategy has already been adopted in Ref. [44] for the Emery model at hole doping. Here we extend it to the electron doped case.

### v2

@@ Abstract
Understanding the similarities and differences between adding or removing electrons from a charge-transfer insulator may provide insights about the origin of the electron-hole asymmetry found in cuprates. Here we study with cellular dynamical mean-field theory the Emery model set in the charge-transfer insulator regime, and dope it with either electrons or holes. We consider the normal state only and focus on the doping evolution of the orbital character of the dopants and on the nature of the doping-driven transition. Regarding the orbital character of the dopants, we found an electron-hole asymmetry: doped electrons mostly enter the copper orbitals, whereas doped holes mostly enter the oxygen orbitals. Regarding the nature of the doping-driven transition, we found no qualitative electron-hole asymmetry: upon either electron or hole doping, there is a two-stage transition from a charge-transfer insulator to a strongly correlated pseudogap and then to a metal. This shows that a strongly correlated pseudogap is an emergent feature of doped correlated insulators in two dimensions, in qualitative agreement with recent experiments on ambipolar Sr1-xLaxCuO2+y cuprate films. Our results indicate that merely doping with holes or electrons a charge-transfer insulator is not sufficient for explaining the electron-hole asymmetry observed in the normal state phase diagram of cuprates. Our work reinforces the view that actual hole-doped cuprates are more correlated than their electron-doped counterparts.
@@ I Introduction
Here, the problem we want to consider is what type of asymmetry results solely from doping a charge-transfer insulator in the normal state with either electrons or holes. First, is there any asymmetry in the doping evolution of the orbital character of the dopant carriers upon electron or hole doping? Second, is there any asymmetry in the qualitative features of the normal state doping-driven transition upon electron or hole doping? To address these questions, we consider the finite-temperature normal state of the Emery model set in the charge-transfer insulating regime and dope the resulting charge-transfer insulator with either electron or holes (see sketch of the density of states of the undoped model in Fig. 1(a)).
The strong correlations giving rise to the charge-transfer insulator present an outstanding theoretical challenge, necessitating the use of nonperturbative methods. Over the years, significant progress on the physics of the Emery model has been enabled by advanced computational techniques, including quantum Monte Carlo methods [20, 21, 22, 23, 24, 25, 26, 27], density matrix renormalization group [28, 29, 30], two-particle self-consistent approach [31, 32], single-site dynamical mean-field theory [33, 34, 35, 36, 37, 8, 17, 38, 39] and its cluster extensions [40, 41, 42, 43, 44, 45, 46, 47, 48], variational cluster approximation [49, 50], density matrix embedding theory [51], and 2D tensor network methods [52]. However, due to the complexity of the Emery model, our understanding of the finite-temperature aspects of the normal state doping-driven metal to charge-transfer insulator transition remain incomplete. Our work employs finite-temperature cellular extension [53, 54, 55] of dynamical mean-field theory [56] (CDMFT) and builds upon Ref. [44], which focused on hole doping only.
@@ II Summary of key findings
Second, we show that the low-temperature normal state doping-driven metal to charge-transfer insulator transition has the same qualitative features upon either electron or hole doping: at zero doping, there is a continuous charge-transfer insulator to strongly correlated pseudogap transition, and at finite doping there is a discontinuous transition between the strongly correlated pseudogap and a more conventional correlated metal (see sequence of phases in Fig. 1(c)). The precise value of doping of the pseudogap to metal transition depends on the dopant carriers, but the qualitative features of the normal state doping-driven transition of a charge-transfer insulator is thus symmetric upon electron or hole doping. This transition is analogous to that found in the doped Mott-Hubbard insulator realised with the single-band 2D Hubbard model [67, 68], as pointed out in Ref. [44] for hole doping.
Our findings that the ambipolar doping of a charge-transfer insulator (i) does produce an asymmetry in the orbital character of the dopants, (ii) but is not sufficient to produce an asymmetry on the qualitative features of the normal state doping-driven transition, enable us to derive two main insights on the microscopic mechanism controlling the doping of a charge-transfer insulator, as we shall discuss in more detail in Section IX.
First, when ambipolar doping of a given charge-transfer insulating system is experimentally attainable, a qualitatively electron-hole symmetry in the normal state doping-driven transition should be expected. Indeed, upon ambipolar doping of the cuprates films of SLCO [9], a surprising symmetry in the high-energy features of the bandstructure of the CuO2 planes is found. Most importantly, our results indicate that a strongly correlated pseudogap, namely a pseudogap solely arising from Mott physics plus short-range correlations, can emerge from a charge-transfer insulator upon either electron or hole doping, i.e. irrespective of the orbital character of the dopant carriers. This finding is in qualitative agreement with the work on ambipolar doping of cuprate films of SLCO [9]. Our finding implies that a strongly correlated pseudogap is an emergent feature of doped correlated insulators in two dimensions.
@@ III.1 Emery model
The Emery model takes into account three hopping processes. An electron can hop from an oxygen orbital (i) to a nearest-neighbor copper orbital via the hopping amplitude $t_{pd}$ (orange arrows), (ii) to a nearest-neighbor (diagonal) oxygen orbital via the hopping amplitude $t_{pp}$ (light blue arrows), and (iii) to an oxygen orbital across the copper orbital via the hopping amplitude $t_{pp}^{\prime}$ (dark blue arrows). The sign of the copper-oxygen and oxygen-oxygen hopping follows the convention indicated in Fig. 2. Let $\epsilon_{d}$ be the energy of the copper $3d_{x^{2}-y^{2}}$ orbitals and $\epsilon_{p}$ be the energy of the oxygen $2p_{x},2p_{y}$ orbitals. Regarding the interaction processes, on a copper orbital two electrons experience a Coulomb repulsion $U_{d}$. For simplicity, we neglect the Coulomb repulsion $U_{p}$ between electrons on an oxygen orbital.
@@ III.2 Cellular dynamical mean-field theory
To solve the Emery model Eq. 1 at finite temperature, we use the cellular extension [53, 54, 55] of dynamical mean-field theory [56] (CDMFT) and follow the procedure described in Ref. [44]. Here we consider the CDMFT equations in the normal state only. The basic idea of CDMFT is to partition the lattice into a superlattice of clusters, to single out one cluster from the lattice and embed it in a self-consistent reservoir of noninteracting electrons. In this work we consider a cluster of $12$ sites with $N_{d}=4$ copper sites and $N_{p}=8$ oxygen sites, as indicated by the green square in Fig. 2.
@@ III.3 Model parameters of this work
Let us also comment our choice of $|\epsilon_{p}-\epsilon_{d}|=8$. This choice follows from the goal of our work to investigate the effect of doping a clear charge-transfer insulator. A smaller value of $|\epsilon_{p}-\epsilon_{d}|$ would be arguably closer to the values obtained in bandstructure calculations for cuprates [70, 8]. However, this would increase the mixed character between the $3d$ and $2p$ orbitals, which, in the Zaanen-Sawatzky-Allen framework [12], moves the Emery model towards the “intermediate” (or covalent or mixed-valence) region. In our work we wish to set the model in a clear charge-transfer insulator regime, that is, away from the intermediate regime. In the language of Ref. [48], our choice of parameters corresponds to an “ionic charge-transfer insulator”, and a smaller value of $|\epsilon_{p}-\epsilon_{d}|$ corresponds to a “covalent charge-transfer insulator”. Furthermore, small values of $|\epsilon_{p}-\epsilon_{d}|$ drastically reduce the Monte Carlo sign, leading to a more severe sign problem.
@@ IV.1 Two-stage metal to charge-transfer insulator transition
To summarise, upon either electron or hole doping, at low temperature the system undergoes a purely electronic two-stage doping-driven metal insulator transition: a continuous charge-transfer insulator to pseudogap transition at $\delta=0$, and a discontinuous pseudpgap to correlated metal transition at finite $\delta$. This two-stage doping-driven metal insulator transition also occurs in the doped Mott insulator realised by the single-band 2D Hubbard model solved with CDMFT [67, 68], as pointed out in Ref. [44].
IV.2 Peak in entropy and enhanced charge compressibility
@@ V Temperature-doping phase diagram
Figure 5: Temperature $T$ vs doping $\delta$ normal state phase diagram of the Emery model solved with CDMFT. Data are for the set of parameters of Eq. 19 and $U_{d}=11.1$. Left: electron doping, right: hole doping. Dark blue squares show the incompressible charge-transfer insulator (CTI) at $\delta=0$. The spinodal line marking the pseudogap (PG) boundary is shown by open red triangles. The spinodal line marking the correlated metal (CM) boundary is shown by open orange triangles. The spinodal lines eventually merge at a critical end point shown by a solid purple circle. The Widom line, determined from the locus of maximum charge compressibility $\kappa$ vs $\delta$ in Fig. 4, extends from the critical end point and is indicated by a dashed purple line with open circles. The cyan dashed line with crosses shows the locus of maximum thermodynamic entropy $s$ vs $\delta$, as calculated by the crossing of the isotherms $\delta(\mu)$. These features occur on both the electron and hole-doped side of the phase diagram but are asymmetrical. Gray filled circles indicate the datapoints studied in our work. Filled hexagons denote the datapoints discussed in Sec. VI.1. Shaded grey area marks the inaccessible region to our calculations due to the Monte Carlo sign problem.
@@
The spinodal lines merge at a second order critical endpoint, for either hole or electron doping. The two filled purple circles are our estimates for the critical end points on the hole and electron doped side of the phase diagram. As a result, the pseudogap to metal transition is a purely electronic transition. Its critical endpoint should belong to the Ising universality class, as shown for the metal to Mott insulator transition [83, 84, 85, 86, 87].
@@
Next, let us compare the hole and electron doping sides of the $T-\delta$ phase diagram. They appear qualitatively similar, but quantitatively different. In both cases, the system shows a first order transition between a pseudogap and a correlated metal which terminates in a critical end point at finite doping and finite temperature, and from which emerges a supercritical crossover characterised by enhanced density fluctuations. By changing the interaction strength $U_{d}$, in Section VIII we will show that the first order transition at hole doping and at electron doping are connected in the $U_{d}-T-\delta$ phase diagram, thereby revealing their common origin in the Mott physics in two dimensions. In other words, in the $U_{d}-T-\delta$ diagram, there is a surface of first-order transitions.
For either electron or hole doping, the pseudogap to metal transition is followed at large doping by a maximum of the thermodynamic entropy $s(\delta)$. The locus of the maximum in $s(\delta)$ at different temperatures, as determined by the zero of the expansion coefficient $(dn_{\rm tot}/dT)_{\mu}$, is shown by a cyan dotted line with crosses: it is only weakly doping dependent and extends up to high temperatures. Physically, the peak in the thermodynamic entropy arises from the localization-delocalization physics of doped Mott insulators, as discussed in Ref. [77, 67, 68, 96, 75, 97] in the context of the single-band 2D Hubbard model. These key features of the thermodynamic entropy $s$ – a doping-independent peak in $s(\delta)$ close to the value of doping where the pseudogap ends – are compatible with experimental results in hole-doped cuprates [78, 79].
Let us now turn to the quantitative differences between the two sides of the $T-\delta$ phase diagram. The temperature and doping position of the critical end point is asymmetric between electron and hole doping. The temperature of the critical end point is higher on the electron-doped side of the phase diagram. The origin of this asymmetry is not evident to us. A similar electron-hole asymmetry in the pseudogap to metal critical endpoint occurs in the single-band 2D Hubbard model on a triangular lattice [98]. This suggests that this asymmetry may be linked to the shape of the Fermi surface upon doping. However further investigations are needed to clarify this aspect. A possible hint may come from the study of the electron-hole asymmetry in the pseudogap to metal critical endpoint in the 2D Hubbard model with next-nearest neighbor hopping on a square lattice [99].
@@
VI Phase characterization: local density of states and spectral function at the Fermi level
@@
The analysis in Sections IV and V reveals the existence of a first-order transition between two compressible metals at finite electron and hole doping. To further characterise the two metallic states separated by the transition and to show that they have the features of a strongly correlated pseudogap and of a correlated metal, we can turn to the study of the local density of states (DOS) and of the spectral function at the Fermi level. The strategy of analyzing the local DOS has already been adopted in Ref. [44] for the hole-doped Emery model. Here we extend it to the electron doped case.
VI.1 Local density of states
@@
VI.2 Spectral function at the Fermi level
Figure 8: Intensity plot of the momentum resolved spectral function at the first Matsubara frequency of the Cu $3d$ orbital ($A_{d}({\bf k}$)) and of the oxygen $2p_{x},2p_{y}$ orbitals ($A_{p}({\bf k})$) in the first quadrant of the Brillouin zone. Data are for the set of parameters of Eq. 19, $U_{d}=11.1$ and $T=1/50$, which lies below the temperature of the pseudogap to metal critical endpoint (see Fig. 5). Panels (a)-(h) correspond to different values of electron doping. Panels (i)-(p) correspond to different values of hole doping.
Figure 8 shows the intensity plot of the spectral function at the first Matsubara frequency $\omega_{0}=i\pi T$ of the Cu $3d$ orbital ($A_{d}({\bf k})=-\frac{1}{\pi}\textrm{Im}G_{d}({\bf k},i\omega_{0})$) and of the O $2p_{x},2p_{y}$ orbitals ($A_{p}({\bf k})=-\frac{1}{\pi}\textrm{Im}G_{p}({\bf k},i\omega_{0})$) at the low temperature $T=1/50$ and for different values of doping. At low temperature, this quantity is a proxy of the spectral function at the Fermi level. Data are obtained for the set of parameters of Eq. 19 and $U_{d}=11.1$. To obtain the translation invariant lattice Green’s functions $G_{d}({\bf k},i\omega_{0})$ and $G_{p}({\bf k},i\omega_{0})$ from the cluster self-energy $\boldsymbol{\Sigma}_{\rm cl}(i\omega_{0})$, we use the Green’s function periodization scheme [101].
Let us first discuss the correlated metal at values of doping larger in absolute value than the first-order transition [third and fourth rows, i.e. panels (e)-(h) for electron doping and panels (m)-(p) for hole doping]. At large $|\delta|$ (fourth row), upon either electron or hole doping, the intensity along the locus of maximum spectral function is large, indicating long-lived excitations, and approximately uniform in ${\bf k}$. Upon electron doping, the overall intensity of $A_{d}({\bf k})$ is larger than that of $A_{p}({\bf k})$. On the other hand, upon hole doping, the overall intensity of $A_{d}({\bf k})$ is comparable to that of $A_{p}({\bf k})$. This reflects the asymmetry regarding which orbital the dopant carriers enter. Upon reducing doping in absolute value [third row], the overall spectral intensity decreases and broadens as a results of the increased correlations.
Next, let us discuss the pseudogap state at small values of either electron or hole doping [first and second rows, i.e. panels (a)-(d) for electron doping and panels (i)-(l) for hole doping]. Upon either electron or hole doping, the intensity along the locus of maximum spectral function is overall small and strongly ${\bf k}$ dependent. This is the hallmark signature of the strongly correlated pseudogap state on the spectral function within cluster DMFT methods, as shown by intensive studies on the single band 2D Hubbard model on a square lattice [69, 102, 103, 104, 105, 106, 107] (for a recent review, see Ref. [108]). Our results extend this finding to the Emery model, and show that a strong ${\bf k}$ dependent depletion of states at the Fermi level occurs at either electron or hole doping.
There is however a quantitative asymmetry between electron and hole doping regarding the region in ${\bf k}$ where the depletion of spectral intensity takes place. For hole doping, the spectral intensity is concentrated along the zone diagonal (i.e., the so-called nodal direction) and is gradually reduced upon approaching the zone boundary (i.e., the so-called antinodal direction). In sharp contrast, for electron doping the spectral intensity is concentrated along the zone boundary and is progressively depleted upon approaching the zone diagonal.
To further characterize the first-order transition between strongly correlated pseudogap and correlated metal, the second and third rows of Fig. 8 show the spectral intensity in the coexistence region: panels (c)-(d) [(k)-(l)] and (e)-(f) [(m)-(n)] have the same chemical potential $\mu$, but different doping levels (see Fig. 3) and different spectral functions. Across the pseudogap to metal first-order transition there is a dramatic depletion and broadening of spectral intensity. The depletion of spectral intensity is especially strong close to the zone diagonal for electron doping, and away from the zone diagonal for hole doping.
Finally, we caution that the calculation of the momentum resolved spectral function in Fig. 8 relies on the periodization – a procedure applied to the converged CDMFT results to extract the translation invariant lattice Green’s function from the cluster self-energy. The periodization is an ill-defined procedure and different periodization schemes have been proposed. Here we used the Green’s function periodization scheme [101]. It has long been debated whether the spectral intensity features are an artefact of the periodization or a result of CDMFT. Recently, Ref. [109] has demonstrated that CDMFT cannot resolve the question of Fermi arcs vs hole pockets, regardless of the periodization used. Nonetheless, the key features of the spectral intensity on the single band 2D Hubbard model solved with CDMFT are compatible with those revealed by other numerical methods [110]. Note also that the phase characterization by the local DOS in Sec. VI.1 does not rely on the periodization procedure.